Question is : Write an equation of the line that is parallel to the graph of 2x-7y=3 and has the same y-intercept as x+3y=9? I keep getting confused, but I think I can always plug in 0 for X to find Y? I think I am doing this wrong. Please help and thank you!
Answer by scott8148(6628)

y-intercept means x=0 ___ (0) + 3y = 9 ___ y = 3
___ so the point is (0,3)

a parallel line will have the same equation EXCEPT for the constant term

substituting ___ 2(0) - 7(3) = ? ___ -21 = ?

the equation is ___ 2x - 7y = -21